OpenStudy (anonymous):

BROS COME HELP..The radius of the Earth is approximately 3,960 miles. Find the approximate surface-area-to-volume ratio of the Earth. A. 0.00025 B. 0.00076 C. 1,320 D. 11,880

OpenStudy (mathstudent55):

You need the formula for the surface of a sphere and the formula for the volume of a sphere.

OpenStudy (anonymous):

@mathstudent55 after i get both what do i do with them?

OpenStudy (anonymous):

Sphere Surface Area = 4 • π • r² = π • d²

OpenStudy (anonymous):

Sphere Volume = 4/3 • π • r³ = ( π •d³)/6

OpenStudy (anonymous):

THANKS BATMAN, YOU'RE MY BRO!

OpenStudy (mathstudent55):

Now you need to divide the surface area by the volume to get the ratio of surface area to volume. ratio \(= \cfrac{1.97061 \times 10^8}{2.6012 \times 10^{11} }\)